OpenStudy (anonymous):

Hi I need some help. Write an equation of the line containing the given point and parellel to the given line. Express your answer in the form y=mx+b (-3,6); 2x=9y+4

OpenStudy (anonymous):

what do parallel lines have in common?

OpenStudy (ash2326):

We have to find a line that's parallel to 2x=9y+4 Let's rewrite this slope intercept form \[2x-4=9y\] \[9y=2x-4\] divide whole equation by 9 \[y=\frac{2}{9} x-\frac{4}{9}\] here slope = 2/9 Parallel Lines' slope are same so our line also has slope =2/9 so \[y=\frac{2}{9}x+c\] now substitute x=-3 and y=6 in this to find C, as this point lies on this line Can you solve it ?

OpenStudy (anonymous):

y=2/9x+c should be if I am substituting -3 and 6 I believe it is like this 6=2/9(-3)+c??